OpenStudy (anonymous):

-8(6w-3v-1) distribute get rid of parenthesis

OpenStudy (sydneeod):

What you would need to do is multiply -8 by everything within the parenthesis. So you have \[-8(6w - 3v - 1)\] First step multiply -8 by the first thing in the parenthesis. That would be 6w. What is 6 times -8 ? remember your variable. then go on to your next number in the parenthesis and so on

jigglypuff314 (jigglypuff314):

Hello :) rather than posting your question over and over in the chat, please just post the link to your question next time the rule works like this a(b+c+d) = ab + ac + ad so -8(6w - 3v - 1) = (-8)(6w) + (-8)(-3v) + (-8)(-1) multiply and simplify :)